# Using iterated local alignment to aggregate trajectory data into a traffic flow map

### v2

@@ Abstract
Desire line maps are widely deployed for traffic flow analysis by virtue of their ease of interpretation and computation. They can be considered to be simplified traffic flow maps, whereas the computational challenges in aggregating small scale traffic flows prevent the wider dissemination of high resolution flow maps. Vehicle trajectories are a promising data source to solve this challenging problem. The solution begins with the alignment (or map matching) of the trajectories to the road network. However even the state-of-the-art map matching implementations produce sub-optimal results with small misalignments. While these misalignments are negligible for large scale flow aggregation in desire line maps, they pose substantial obstacles for small scale flow aggregation in high resolution maps. To remove these remaining misalignments, we introduce innovative local alignment algorithms, where we infer road segments to serve as local reference segments, and proceed to align nearby road segments to them. With each local alignment iteration, the misalignments of the trajectories with each other and with the road network are reduced, and so converge closer to a minimal flow map. By analysing a set of empirical trajectories collected in Hannover, Germany, we confirm that our minimal flow map has high levels of spatial resolution, accuracy and coverage.
Keywords: Desire lines, floating car data FCD, GPS, map matching, route finding
@@ 1 Introduction
One of the fundamental quantities in transport planning is a traffic flow map, i.e. a map of the traffic flow levels on the road segments in a road network (Ortúzar and Willumsen, 2011). Whilst traffic flow maps are a rich source of information about vehicle mobility patterns, they are costly in terms of time and resources to compute for any reasonably sized road network. To alleviate this cost burden, most approaches restrict the spatial coverage and resolution of the flow map. One of the most common is to place sensors at fixed locations in the road network, whose results are then visualised as a traffic count map. Another are trip intent/recall questionnaires to inform large scale properties such as trajectory origins and destinations, which can be visualised as a desire line flow map/spider diagram (Tobler, 1987). Both of these are simplified flow maps, since the detailed mobility patterns outside of the sensor locations or origin/destination pairs are not known. These unknown patterns can be inferred from other data sources, such as route assignment models (Evans, 1976). While these model-assigned routes are highly detailed, the trade-off is that they are not guaranteed to correspond closely to empirical mobility patterns.
@@
Thus an ideal data source combines the empirical information of road sensor counts or questionnaires, with the small scale details of model-assigned trajectories. This gap in the market can be filled by trajectory data. Due to the prevalence of geolocation-enabled devices, such as vehicle navigation guides and mobile telephones, trajectory data can be acquired with low marginal cost whilst at the same time offering extensive spatial coverage and resolution of empirical mobility patterns (Herrera et al., 2010, Andrienko and Andrienko, 2013). Throughout, we employ the term ‘trajectory data’ rather than Floating Car Data (FCD) or Floating Mobile Data (FMD), as our analysis is not restricted to trajectory data from cars or from mobile phones, so these terms are equivalent for our purposes. An example is the 1183 trajectories are collected from a GPS-enabled mobile phone, from December 2017 to March 2019 by a single driver in Hannover, Germany, with an overall average sampling rate about 1 point per second (Zourlidou et al., 2022). They are plotted as the green circles in Figure 1.
Refer to caption(a)
Refer to caption(b)
Refer to caption(c)
Figure 1: Trajectories in Hannover, Germany. (a) City level. (b) Neighbourhood level, zoom of black rectangle. (c) Small neighbourhood level, zoom of blue rectangle. Trajectory ID = 7 (orange circles), 315 (purple diamonds), others (green circles).
In Figure 1(a), at the city level, the points (green circles) appear to be aligned to the road network. If we zoom in on the small black rectangle in the central region, then in Figure 1(b) at the neighbourhood level, we observe that the points deviate from the road network. This deviation is clearer in the closer zoom in Figure 1(c). Moreover, if we focus on the orange circles (Trajectory ID = 7) and purple diamonds (ID = 315), then we observe that the vehicle location between the recorded points is unknown. These maps illustrate the errors in trajectories. These are broadly classified as ‘measurement errors’ where the recorded coordinates are not the true locations, and ‘sampling errors’ where the information about the trajectory is lost in between recorded coordinates (Ortúzar and Willumsen, 2011, Saki and Hagen, 2022).
Our goal is to produce a traffic flow map from these noisy trajectories, which can be utilised at any scale, from the city/regional level to the individual road segment level. This requires us to minimise the errors in trajectories. Our approach is composed of two stages. The first stage is to align the trajectories to the road network, which is known as map matching (Quddus et al., 2007, Chao et al., 2020). It produces a route, which is a connected sequence of road segments in the road network, that is consistent with the trajectory. Our contribution is an improvement to standard map matching by adding post hoc route finding. In common with many open source transport planning tools, we employ APIs based on the OpenStreetMap (OSM) network. While we are able to improvement the overall alignment to the road network, these map matched routes inherit incompressible misalignments, ranging from several centimetres to several metres, from the OSM road network. These small misalignments prevent the accurate aggregation of traffic flows at this scale.
@@
The second stage is to resolve these misalignments, and our contribution here is the proposed local alignment of map matched routes. In contrast to global alignment approaches, we locally infer which road segments should serve as a local reference, and then proceed to align other nearby road segments to it. To accomplish this, we introduce several novel algorithms which employ a mix of advanced statistical and geospatial methods. These include ‘node snapping’ where the nearby boundary points of road segments are combined via statistical clustering, and ‘line blending’ where road segments, which are near to each other but do not share overlapping road sub-segments, are aligned to maximise their overlapping sub-segments. Inputting these locally aligned routes into a flow aggregation API leads to a more accurate flow map. Iterating these local alignments in turn leads to a minimal flow map.
The outline of the paper is as follows. In Section 2, we describe the computation of map matched routes using off-the-shelf map matching and route finding APIs. In Section 3, we describe the local alignment of the map matched routes and their aggregation into a flow map. In Section 4, we demonstrate that the locally aligned flow map from the Hannover trajectories is well-aligned to the OSM road network, and has a high level of accuracy and spatial coverage of estimated traffic flows compared to reference traffic flows. We then discuss some software implementation issues, and some concluding remarks.
2 Alignment of trajectories with map matching and route finding
@@
We introduce some mathematical notation to state precisely the problem of map matching. We represent the road network by a graph $\mathcal{N}=\mathcal{N}(\mathcal{E},\mathcal{V})$, where the $n_{\mathcal{E}}$ edges =$\mathcal{E}=\{{\bm{e}}_{1},\dots,{\bm{e}}_{n_{\mathcal{E}}}\}$ of this graph are the road segments and the $n_{\mathcal{V}}$ nodes/vertices $\mathcal{V}=\{{\bm{v}}_{1},\dots,{\bm{v}}_{n_{\mathcal{V}}}\}$ indicate that two (or more) different road segments are accessible to/from each other at this node point. These nodes ${\bm{v}}_{i}$ are single points. Each road segment is composed of a sequence of connected piecewise linear segments, so ${\bm{e}}_{i}=\{{\bm{e}}_{i,1},\dots,{\bm{e}}_{i,n_{i}}\}$ is an ordered sequence of $n_{i}$ points ${\bm{e}}_{i,j},j=1,\dots,n_{i}$. We refer to this as a ‘linestring’ geometry, following the Open Geospatial Consortium terminology (OGC, 2010). We set $\mathcal{N}$ to be the OSM network which is a freely available road network with global coverage.
We denote a single trajectory $G=\{{\bm{g}}_{1},\dots,{\bm{g}}_{n_{G}}\}$ as a temporally ordered sequence of $n_{G}$ points ${\bm{g}}_{i},i=1,\dots,n_{G}$. This is known as a ‘multipoint’ geometry (OGC, 2010). Whilst some authors require that each point ${\bm{g}}_{i}$ be accompanied by their timestamp $t_{i}$ to be considered a trajectory, this is not strictly required since ${\bm{g}}_{i}$ are ordered according to the timestamps, even if the timestamps themselves are not recorded in the trajectory. Due to measurement error, the points of a trajectory $G$ are not necessarily coincident with the road network $\mathcal{N}$, and due to sampling error, there is no information about the vehicle in between the points of $G$.
We represent the output of a map matching algorithm $M$ from an empirical trajectory $G$ as $M(G)=\{{\bm{m}}_{1},\dots,{\bm{m}}_{n_{M}}\}$ which is an ordered, connected sequence of $n_{M}$ edges. The goal is that the map matched route follows closely the road network $\mathcal{N}(\mathcal{V},\mathcal{E})$. Whilst it is straightforward to ensure that all boundary points of the segments $\{{\bm{m}}_{1},\dots,{\bm{m}}_{n_{M}}\}$ coincide with the nodes $\mathcal{V}$ of the road network graph $\mathcal{N}$, it is more challenging to ensure that the segments $\{{\bm{m}}_{1},\dots,{\bm{m}}_{n_{M}}\}$ themselves coincide with the edges $\mathcal{E}$. In comparison to an empirical trajectory $G$, all boundary points of the edges of a map matched route $M(G)$ are aligned to the road network (reduced measurement error) and the vehicle position is estimated by the linestring connecting the boundary points of the road segment (reduced sampling error).
@@
There is a large body of research on this difficult problem of map matching. We focus on the popular class of Hidden Markov Models (HMM) map matching algorithms. HMM methods iteratively build the map matched route by selecting the most likely next segment to connect to the current route using a probabilistic model. According to a review of map matching algorithms (Chao et al., 2020), HMM is a state-transition method. The other three classes are similarity, candidate-evolving and scoring methods. Further details of alternative map matching algorithms are found in Quddus et al. (2007), Chao et al. (2020). We leave this discussion here since the improvements offered by our proposed methods are valid for any map matching algorithm, and concentrate on HMM map matching due to its accuracy and computational efficiency. Even if we restrict ourselves to HMM algorithms on the OSM road network, there are many off-the-shelf map matching APIs available. We focus on the Valhalla routing engine (https://valhalla.github.io/valhalla), which includes its highly recommended map matching API (Saki and Hagen, 2022).
Figure 2 displays the $n=1147$ map matched routes by the Valhalla map matching API. We discard 36 trajectories (3.04%) from our original data set of 1183 trajectories. In Figure 2(a), the map matched routes (blue lines) overall are well-aligned to the road network. For the orange map matched route, all its segments are aligned to the road network, whereas the purple route appears to be displaced by several metres from the road centreline. The measurement and sampling errors of the map matched routes $M(G)$ are reduced in comparison those for the empirical trajectories $G$, though these errors remain sizeable at the road segment level in Figure 2(b).
Refer to caption(a) Refer to caption(b)
Figure 2: Map matched routes, using only Valhalla map matching API. (a) Neighbourhood level. (b) Small neighbourhood level, zoom of black rectangle. Trajectory ID = 7 (orange), 315 (purple), others (blue).
@@
Our first contribution is to better align the edges of the map matched routes $M(G)$ to the road network edges $\mathcal{E}$. We propose a post hoc adjustment of the map matching output by an additional call to a route finding API, as outlined in Algorithm 1. The inputs of ST_ROUTE are the empirical trajectory $G$, the map matching API $M$, the route finding API $R$, and the number of waypoints ${\bm{n}}_{W}$ for the route finding. Since we do not have an a priori single optimal value for the number of waypoints, we consider a range of $w$ values ${\bm{n}}_{W}=(n_{W,1},\dots,n_{W,w})$. In Step 1, we compute the initial map matched route $M(G)$ from the empirical trajectory $G$ by calling the map matching API $M$. This initial map matched route $M(G)=\{{\bm{m}}_{1},\dots,{\bm{m}}_{n_{M}}\}$ is a linestring with $n_{M}$ edges, with $n_{M}+1$ points. In Steps 2–5, we loop over the $w$ number of waypoints in ${\bm{n}}_{W}$. In Steps 3–4, we take a sample of $n_{W,i}$ waypoints, where the first waypoint ${\bm{w}}_{1}$ is the start point of ${\bm{m}}_{1}$ and the $n_{W,i}$th way point is the end point of ${\bm{m}}_{n_{M}}$, and the intermediate waypoints ${\bm{w}}_{2},\dots,{\bm{w}}_{n_{W,i}-1}$ are sampled from the start points of the edges ${\bm{m}}_{2},\dots,{\bm{m}}_{n_{M}}$. In Step 5, we call the route finding API $R$ with the waypoints ${\bm{w}}_{1},\dots,{\bm{w}}_{n_{W,i}}$. The result is the map matched route $M_{i}^{*}(G)=R({\bm{w}}_{1},\dots,{\bm{w}}_{n_{W,i}})=\{{\bm{m}}_{1}^{*},\dots,{\bm{m}}_{n_{M^{*}}}\}$, which is a route of $n_{M^{*}}$ connected edges. In Step 6, we select the route with the smallest dynamic time warping (DTW) normalised distance between the routes $M_{i}^{*}(G)$ and the empirical trajectory $G$, for $i=1,\dots w$. The DTW distance is based on the lengths of all the distortions of $M_{i}^{*}(G)$ to achieve a maximal alignment between $M_{i}^{*}(G)$ and $G$ (Sakoe and Chiba, 1978, Giorgino, 2009).
@@
2:Output: $M^{*}(G)$ map matched route
@@
6:     Sample $n_{W}$ waypoints ${\bm{w}}_{1},\dots,{\bm{w}}_{n_{W}}$ from $M(G)$, with ${\bm{w}}_{1}$ := $\operatorname{Start}(M)$, ${\bm{w}}_{n_{W}}:=\operatorname{End}(M)$
7:     Compute map matched route $M^{*}_{i}(G):=R({\bm{w}}_{1},\dots,{\bm{w}}_{n_{W}})$
8:Select minimal route $M^{*}(G):=\operatorname{argmin}_{i\in\{1\dots w\}}\,\operatorname{DTW}(M^{*}_{i}(G),G)$
We set $R$ to be the Valhalla Odin turn-by-turn route finding API to be consistent with our choice of $M$ as the Valhalla Meili map matching API. We set the number of waypoints as ${\bm{n}}_{W}=3,13,23,33,43,63,83$. Figure 3 displays the results $M^{*}(G_{1}),\dots,M^{*}(G_{1147})$ from ST_ROUTE. For the trajectory ID = 7 (orange) $n_{W}=83$, and ID = 315 (purple) $n_{W}=23$ give the minimal DTW route. The overall impression of the map matched routes $M^{*}$ in Figure 3(a) is that the misalignment has been reduced, especially in the purple line since it now aligns more closely to the road centreline. In Figure 3(b), at the level of road segments, whilst the map matched routes $M^{*}$ tend to be contained inside the road segments, they are not exactly coincident with each other. This is in part because the road network graph $\mathcal{N}$ itself contains small misalignments, and so they are propagated into any map matching or route finding algorithm based on it. These small alignments lead to a lack of overlapping sub-segments, which in turn lead to inaccurate flow aggregation.
Refer to caption(a) Refer to caption(b)
Figure 3: Map matched routes, using Valhalla map matching and route finding APIs. (a) Neighbourhood level. (b) Small neighbourhood level, zoom of black rectangle. Trajectory ID = 7 (orange), 315 (purple), others (blue).
3 Local alignment of road segments for flow map aggregation
Our goal is to resolve the remaining misalignments from the map matching/route finding in the previous section, so we are able to aggregate accurately traffic flows on road segments. We aim to achieve this by local, internal alignment between the map matched routes. By internal alignment, we mean that we align the routes with each other, rather than to the external road network graph. Since an external reference road network is not required as an explicit input, our proposal can be deployed in more cases, e.g. when the quality of the road network graph is insufficient, or when the alignment to the road network graph is computationally intensive. By local alignment, we mean that we focus on aligning sub-segments of the routes, rather than complete routes.
[…]
We begin by illustrating the difference in flow aggregation between the $M$ and $M^{*}$ map matched routes. For the routes $M$ from Figure 2(b) and $M^{*}$ routes from Figure 3(b), the flow maps are given below in Figure 4(a–b) respectively. In these maps, the colour (purple to orange) and width of the road segments is proportional to the traffic flow. We observe that there are fewer, wider linestring segments in Figure 4(b) than in Figure 4(a).
Refer to caption(a) Refer to caption(b)
Figure 4: Flow maps for map matched routes. (a) Map matched routes with Valhalla map matching only. (b) Map matched routes with Valhalla map matching and route finding. Colour (purple to orange) and width of road segments is proportional to traffic flow.
The flow aggregation in Figure 4 was carried out using the overline function in the R package stplanr, which we refer to as ST_OVERLINE_PLANR (Lovelace and Ellison, 2018). Starting with the map matched routes =$\mathcal{M}^{*}=\{M^{*}({G_{1}}),\dots,M^{*}(G_{n})\}$, then the flow map is $\mathcal{F}={\tt ST\_OVERLINE\_PLANR}(\mathcal{M}^{*})$. This flow aggregation involves the search for all road segments from the routes which exactly equal to each other. These exactly equal road segments are reduced to a single common segment, and the associated traffic flow is the number of exactly equal segments. Since it relies on exactly equal road segments, then small misalignments are sufficient to make the flow aggregations inaccurate.

### v5

@@ Abstract
Vehicle trajectories are a promising GNSS (Global Navigation Satellite System) data source to compute multi-scale traffic flow maps ranging from the city/regional level to the road level. The main obstacle is that trajectory data are prone to measurement noise. While this is negligible for city level, large-scale flow aggregation, it poses substantial difficulties for road level, small-scale aggregation. To overcome these difficulties, we introduce innovative local alignment algorithms, where we infer road segments to serve as local reference segments, and proceed to align nearby road segments to them. We deploy these algorithms in an iterative workflow to compute locally aligned flow maps. By applying this workflow to synthetic and empirical trajectories, we verify that our locally aligned flow maps provide high levels of accuracy and spatial resolution of flow aggregation at multiple scales for static and interactive maps.
Keywords: GNSS, GPS, rasterisation, multi-scale, route finding
@@ 1 Introduction
One of the fundamental quantities in transport planning is a traffic flow map, i.e., a map of the traffic flow levels on the road segments in a road network (Ortúzar and Willumsen, 2011). Whilst traffic flow maps are a rich source of information about vehicle mobility patterns, they are costly in terms of time and resources to compute for any reasonably sized road network. To alleviate this cost burden, most approaches restrict the spatial resolution of the flow map. One of the most common is to place sensors at fixed locations in the road network, whose results are then visualized as a traffic count map. Another common approach are trip intent/recall questionnaires to inform large-scale properties such as trajectory origins and destinations, whose results are then visualized as a desire line flow map/spider diagram (Tobler, 1987). Both of these are simplified flow maps, since the detailed mobility patterns outside of the sensor locations or origin/destination pairs are not known. These unknown patterns can be inferred from other data sources, such as route assignment models (Evans, 1976). While these model-assigned routes are highly detailed, the trade-off is that they are not guaranteed to correspond closely to empirical mobility patterns.
@@
Trajectory data, derived from GNSS (Global Navigation Satellite Systems), combine the large-scale information of road sensor counts or questionnaires, with the small-scale details of model-assigned trajectories. Due to the prevalence of GNSS-enabled devices, such as vehicle navigation guides and mobile telephones, trajectory data can be acquired with low marginal cost whilst at the same time offering extensive spatial coverage and resolution of empirical mobility patterns (Herrera et al., 2010, Andrienko and Andrienko, 2013). Our motivating example of trajectory data are the 1183 trajectories in EPSG 32632 (WGS 84/UTM zone 32N) projected coordinates, collected from a GNSS-enabled mobile phone, from December 2017 to March 2019 by a single driver in Hannover, Germany, with an overall average sampling rate about 1 point per second (Zourlidou et al., 2022). In Figure 1(a), at the city level, the green circles appear to be aligned to the road network.
Figure 1: Trajectory data in Hannover, Germany. (a) Empirical trajectories (green circles). Klusreide road segment (solid yellow circle). (b) Zoom of trajectories and map matched routes (blue lines), with road network (solid black lines), near Klusreide. (d) Zoom of 2 m $\times$ 2 m rasterised flow aggregation. Colour scale is traffic flow in grid cells.
If we zoom in on these trajectory points around Klusriede (solid yellow circle), a residential street near the A2 autobahn, in Figure 1(b), then we observe that the points deviate from the road network (solid black lines). These noisy trajectories are unsuitable for direct aggregation into a flow map, so the usual first step is to apply a map matching procedure (Quddus et al., 2007, Chao et al., 2020). The map matched routes are the blue lines in Figure 1(b). These follow the road network more closely. Map matching is a highly difficult problem and map matching with exact alignment to the underlying road network in a reasonable time frame is usually not feasible. The approach for fast computations taken by the Valhalla routing engine introduces some approximations which do not lead to exact alignment of empirical trajectories with the road network or to each other (Sghaier, 2019). Nonetheless the misalignment is reduced to less than 5 m, which is, for many purposes, including most notably navigation, indeed a suitable result.
For these inexactly aligned trajectories, applying the state-of-the-art traffic flow aggregation API (Morgan and Lovelace, 2021) leads to an inaccurate flow map. The authors resolve this by rasterising the flow map to a raster grid: and within each raster grid cell, the flow is the sum of the flows of all trajectories which intersect the grid cell. This rasterised flow map is displayed in Figure 1(c). The colour scale for the flow ranges from 100 (violet) to 300 (pink) to 600 (orange). The 2 m $\times$ 2 m raster grid obscures the road network at this scale. The resolution of a rasterised aggregation flow map is fixed by the resolution of the raster grid. If we require a higher resolution, then we must recalculate the flow map on a raster grid with a higher resolution.
@@
Our goal is to produce a traffic flow map from map matched routes, which can be utilised at any scale without recalculation, from the city level to the individual road level. In our proposed local alignment approach, we infer which road segments should serve as a local reference, and then proceed to align other nearby road segments to it. To accomplish this, we introduce several novel algorithms which employ a mix of advanced statistical and geospatial methods. These include “node snapping” where the nearby boundary points of road segments are combined via statistical clustering, and “line blending” where road segments, which are near each other but do not share overlapping sub-segments, are aligned to maximise their overlapping sub-segments. We combine these novel algorithms into an iterative workflow which outputs a locally aligned flow map. This flow map is composed of road segments rather than raster grid cells, so its resolution is not limited by the raster grid resolution.
The outline of the paper is as follows. In Section 2, we construct an iterative workflow which computes a locally aligned flow map from noisy trajectories. In Section 3, we apply this workflow to synthetic trajectories from Los Angeles, USA and empirical trajectories from Hannover, Germany. We verify that, with a suitable tuning parameter choice, the locally aligned flow maps have high resolution and spatial accuracy. We then discuss some computational issues and end with some concluding remarks.
2 Local alignment of route segments for flow aggregation
@@
We represent the road network by a graph $\mathcal{N}=\mathcal{N}(\mathcal{E},\mathcal{V})$, where the $n_{\mathcal{E}}$ edges =$\mathcal{E}=\{{\bm{e}}_{1},\dots,{\bm{e}}_{n_{\mathcal{E}}}\}$ are the road segments and the $n_{\mathcal{V}}$ nodes/vertices $\mathcal{V}=\{{\bm{v}}_{1},\dots,{\bm{v}}_{n_{\mathcal{V}}}\}$ indicate that different road segments are accessible to/from each other at this node point. We set $\mathcal{N}$ to be the OpenStreetMap (OSM) road network. Each path in the network is composed of a connected sequence of edges, and we refer to this as a “linestring” geometry, following the Open Geospatial Consortium (OGC) terminology (OGC, 2010). We denote a single trajectory $G=\{{\bm{g}}_{1},\dots,{\bm{g}}_{n_{G}}\}$ as a temporally ordered sequence of $n_{G}$ points ${\bm{g}}_{i},i=1,\dots,n_{G}$. This is a “multipoint” geometry (OGC, 2010). The points of a trajectory $G$ are not necessarily coincident with the road network $\mathcal{N}$.
2.1 Map matching
@@
We represent the output of a map matching algorithm $M$ from a trajectory $G$ as $M(G)=\{{\bm{m}}_{1},\dots,{\bm{m}}_{n_{M}}\}$ which is a connected sequence of $n_{M}$ edges, i.e. a linestring geometry. In comparison to a trajectory $G$, all boundary points of the edges of a map matched route $M(G)$ are aligned to the road network. There is a large body of research on this difficult problem of map matching. We focus on the popular class of Hidden Markov Models (HMM) map matching algorithms. HMM methods iteratively build the map matched route by selecting the most likely next segment to connect to the current route using a probabilistic model. According to a review of map matching algorithms (Chao et al., 2020), HMM is a state-transition method. The other three classes are similarity, candidate-evolving, and scoring methods. Further details of these alternatives are found in Quddus et al. (2007), Chao et al. (2020). We leave this discussion here since our proposed methods are valid for any map matching algorithm, and concentrate on HMM map matching due to its accuracy and computational efficiency. Even if we restrict ourselves to HMM algorithms on the OSM road network, there are many off-the-shelf map matching APIs available. We focus on the Valhalla routing engine (https://valhalla.github.io/valhalla), which includes its highly recommended map matching API (Saki and Hagen, 2022).
@@
To this off-the-shelf Valhalla map matching, we apply a post hoc modification to improve its suitability as input for flow map aggregation, as outlined in Algorithm 1. The inputs of ST_ROUTE are the trajectory $G$, the map matching API $M$, the route finding API $R$, and the number of waypoints ${\bm{n}}_{W}$ for the route finding. These waypoints guide the route finding API to pass near them. Since we do not have an a priori single optimal value for the number of waypoints, we consider a range of $w$ values ${\bm{n}}_{W}=(n_{W,1},\dots,n_{W,w})$. In Step 1, we compute the initial map matched route $M(G)$ from the trajectory $G$ by calling the map matching API $M$. This initial map matched route $M(G)=\{{\bm{m}}_{1},\dots,{\bm{m}}_{n_{M}}\}$ is a linestring with $n_{M}$ edges, with $n_{M}+1$ points. In Steps 2–5, we loop over the number of waypoints in ${\bm{n}}_{W}$. In Steps 3–4, we take a sample of $n_{W,i}$ waypoints, where the first waypoint ${\bm{w}}_{1}$ is the start point of ${\bm{m}}_{1}$ and the $n_{W,i}$th waypoint is the end point of ${\bm{m}}_{n_{M}}$, and the intermediate waypoints ${\bm{w}}_{2},\dots,{\bm{w}}_{n_{W,i}-1}$ are sampled from the start points of the edges ${\bm{m}}_{2},\dots,{\bm{m}}_{n_{M}}$. In Step 5, we call the route finding API $R$ with the waypoints ${\bm{w}}_{1},\dots,{\bm{w}}_{n_{W,i}}$. The result is the map matched route $R_{i}(G)=R({\bm{w}}_{1},\dots,{\bm{w}}_{n_{W,i}})$. In Step 6, we select the route with the smallest dynamic time warping (DTW) normalised distance between the routes $R_{i}(G)$ and the trajectory $G$, for $i=1,\dots w$. The DTW distance is based on the lengths of all the distortions of $R_{i}(G)$ to achieve maximal alignment between $R_{i}(G)$ and $G$ (Sakoe and Chiba, 1978, Giorgino, 2009).
@@
2:Output: $R(G)$ map matched route
@@
6:     Sample waypoints ${\bm{w}}_{1},\dots,{\bm{w}}_{n_{W}}$ from initial map matched $M(G)$
7:     Compute map matched routes $R_{i}(G):=R({\bm{w}}_{1},\dots,{\bm{w}}_{n_{W}})$
8:Select minimal route $R(G):=\operatorname{argmin}_{i\in\{1,\dots,w\}}\,\operatorname{DTW}(R_{i}(G),G)$
We set $R$ to be the Valhalla Odin turn-by-turn route finding API to be consistent with our choice of $M$ as the Valhalla Meili map matching API. We set the number of waypoints as ${\bm{n}}_{W}=3,13,23,33,43,63,83$ since these ${\bm{n}}_{W}$ cover a range of possible values which likely includes an approximately optimal value. The map matched Hannover trajectories are displayed as the blue lines in Figure 1.
2.2 Flow aggregation
To the road network graph $\mathcal{N}=\mathcal{N}(\mathcal{E},\mathcal{V})$, we add the traffic flows on each of the network edges to form a flow map =$\mathcal{F}=\{(f,{\bm{f}}):{\bm{f}}\in\mathcal{N},f>0\}=\{(f_{1},{\bm{f}}_{1}),\dots,(f_{n_{\mathcal{F}}},{\bm{f}}_{n_{\mathcal{F}}})\}$ composed of $n_{\mathcal{F}}$ flows, where each ${\bm{f}}_{i}$ is composed of edges in $\mathcal{E}$, with traffic flow $f_{i}$, $i=1,\dots,n_{\mathcal{F}}$. Our goal is to estimate the traffic flows $f_{1},\dots,f_{n_{\mathcal{F}}}$ and the flow network graph $\mathcal{F}$. For exactly aligned trajectories, to compute the flow aggregation, we deploy the overline function in the R package stplanr, which we refer to as ST_OVERLINE_PLANR (Lovelace and Ellison, 2018). That is, starting with the map matched routes =$\mathcal{R}=\{R(G_{1}),\dots,R(G_{n})\}$, the flow map is $\mathcal{F}={\tt ST\_OVERLINE\_PLANR}(\mathcal{R})$.